Marine, Inc., manufactures a product that is available in both a flexible and a rigid model. The company has made the rigid mActivity Cost Pool and Activity Measure Purchase orders (number of orders) Rework requests (number of requests) Product testiComplete this question by entering your answers in the tabs below. Req 1A Req 1B Req 1B | Req 2 Rega Req ЗА Req 3A Reg 3B ReqComplete this question by entering your answers in the tabs below. Req 1A Req 1B Reg 2 Req 3A Req 3B ------- Compute the actiComplete this question by entering your answers in the tabs below. Req 1A Req 1B Req 2 Reqz | Req3A | Req 3A Req 3B Using actComplete this question by entering your answers in the tabs below. Req 1A Req 1B Reg 2 Req ЗА Req 3B Using activity-based cos
Answer #1

Solution 1a:

Total direct labor hours = (1000*2) + (10000*1) = 12,000

Predetermined overhead rate = $600000/ 12000 = $50 per DLH

Solution 1b:

Computation of Unit Product Cost
Flexible Rigid
Direct Materials $110.00 $80.00
Direct Labor $30.00 $15.00
Overhead Cost per unit ($50*Labor hours per unit) $100.00 $50.00
Unit product cost $240.00 $145.00

Solution 2:

Computation of Activity rates
Activity cost pool Overhead costs Estimated activity Activity Rate
Purchase orders $20,000 400 $50.00 per order
Rework requests $10,000 200 $50.00 per request
Product testing $2,10,000 2100 $100.00 per test
Machine Related $3,60,000 4000 $90.00 per MH

Solution 3a and 3b:

Computation of Overhead allocated and Overhead per unit
Flexible Rigid
Activity cost pool Activity Rate Estimated Activity Overhead Allocated Estimated Activity Overhead Allocated
Purcahse orders $50.00 100 $5,000 300 $15,000
Rework requests $50.00 60 $3,000 140 $7,000
Product testing $100.00 900 $90,000 1200 $1,20,000
Machine Related $90.00 1500 $1,35,000 2500 $2,25,000
Total Overhead allocated $2,33,000 $3,67,000
No. of units 1000 10000
Overhead per unit 233.00 36.70
Computation of Unit Product Cost
Flexible Rigid
Direct Materials $110.00 $80.00
Direct Labor $30.00 $15.00
Overhead Cost per unit $233.00 $36.70
Unit product cost $373.00 $131.70